OpenStudy (anonymous):

4.3 x 10^-3/8.6 x 10^-7 Evaluate and write in scientific notation

OpenStudy (valpey):

\[\frac{4.3 * 10^{-3}}{8.6 * 10^{-7}} = \frac{4.3}{8.6}*10^{-3+7} = \frac{1}{2}*10^{4} = \frac{10}{2}*10^{3} = 5*10^{3} \]

OpenStudy (valpey):

I mean: \[5.0*10^{3}\]

OpenStudy (valpey):

The 5.0 is actually kind of important.